OpenStudy (anonymous):

what fraction of 2 hours is 24 minutes?

OpenStudy (anonymous):

(60+60)/24 (Don't divide. It's a fraction) Simplify it.

OpenStudy (anonymous):

so u change the hour to minutes, u get 120 mins fro 2 hours therefore, 24/120 = 1/5

OpenStudy (anonymous):

@listongeorge r u there?

OpenStudy (anonymous):

You can also do it this way 2h24min = 144 min 1h --- 60min x --- 144min 60x = 144 x = 144/60 simplifying: x = 72/30 x = 36/15

OpenStudy (anonymous):

@listongeorge who r u chanting to?